OpenStudy (chrisplusian):

a particles position on the x axis is given by the function x=(t^2 -4t +2)m where t is is in s. where is the particle when v=4.0 (m/s)????

OpenStudy (chrisplusian):

so what I did was differentiate the function and got the velocity function 2t-4

OpenStudy (chrisplusian):

then I set that equal to 4 "2t-4=4" and solved for t. That gave me t=4. Then I tried to plug that back into the position function and it gave me x=10m. but that answer didn't check with the solution guide, which said x=2m. What did I do wrong?

OpenStudy (anonymous):

plugging t=4 back into \(\large t^2-4t+2 \) gives you 2..... \(\large 4^2-4(4)+2=16-16+2=2 \)

OpenStudy (chrisplusian):

woops! THANKS THAT WAS A BAD MOMENT!!!